Phosphoric acid,  H3PO4,H3PO4, has some important applications. It can be used to produce fertilizers and it is present in soft drinks. Phosphoric acid can be made from phosphorus in a two step process:

Reaction 1: P4+5O2→P4O10P4+5O2→P4O10   

Reaction 2:  P4O10+6H2O→4H3PO4P4O10+6H2O→4H3PO4  

  1. What is the molar mass of phosphoric acid,  H3PO4?H3PO4?  Report your answer to 3 significant figures.

g/mol

  1. How many moles of H3PO4H3PO4 can be produced from one mole of  P4P4 ? Hint: You need to look at both reaction (1) and reaction (2). Your answer should be a simple whole number

mol

  1. A synthesis yielded 235 g of H3PO4.H3PO4. How many moles of H3PO4H3PO4  were produced?

mol

  1. How many moles of P4P4  would be required to produce the number of moles of H3PO4H3PO4  determined in question 3?

mol

  1. Convert the moles of P4P4  determined in question 4 into grams. This will tell us the grams of  P4P4  that actually reacted. Hint: Here you need to use the molar mass of P4,P4, not the molar mass of P.

g

  1. Let's say that the experimentalist actually used 175 g from a bottle labeled  P4.P4. Based on your answer to question 5, you can see that it did not all react. We will assume that this was because the reagent in this bottle contained inert impurities. Based upon your answer to question 5 and the mass of the reagent taken from the bottle labeled  P4,P4, what was the actual mass percent of P4P4  in the bottle? Hint: Be sure to convert into a percentage (by multiplying the fraction by 100%).

%w/w
Answer #1

P4 + 5 O2\rightarrow P4O10

P4O10 + 6 H2O \rightarrow 4 H3PO4

Q1:

Molar mass of H3PO4 = 3×1.008 + 30.9738 + 4×15.9994 = 97.9952g/mol = 98.0 g/mol. (Answer)

Q2:

1 mole of P4 would produces 4 moles of phosphoric acid .

4 mol. (Answer)

Q3:

Number of moles of phosphoric acid produced = 235g/97.9952g/mol = 2.398mol = 2.40 mol. (Answer)

Q4:

The number of moles of P4 needed = moles of phosphoric acid /4 = 2.398mol /4 = 0.5995mol = 0.600 mol. (Answer)

Q5:

The mass of P4 needed = 0.5995mol×123.895g/mol = 74.277g = 74.3 g (Answer)

Q6:

Mass percent of P4 in the bottle = (74.277g/175g)×100%

= 42.444 %

= 42.4 %. (Answer)